# Neuro-Symbolic Integration Brings Causal and Reliable Reasoning Proofs

### v2

@@ Abstract
Two lines of approaches are adopted for complex reasoning with LLMs. One line of work prompts LLMs with various reasoning structures, while the structural outputs can be naturally regarded as intermediate reasoning steps. Another line of work adopt LLM-free declarative solvers to do the reasoning task, rendering higher reasoning accuracy but lacking interpretability due to the black-box nature of the solvers. Aiming to resolve the trade-off between answer accuracy and interpretability, we present a simple extension to the latter line of work. Specifically, we showcase that the intermediate search logs generated by Prolog interpreters can be accessed and interpreted into human-readable reasoning proofs. As long as LLMs correctly translate problem descriptions into Prolog representations, the corresponding reasoning proofs are ensured to be causal and reliable. On two logical reasoning and one arithmetic reasoning datasets, our framework obtains significant improvements in terms of both answer accuracy and reasoning proof accuracy. We released our code at https://anonymous.4open.science/r/CaRing-477B for future research regarding better reasoning proofs using LLMs.
@@
Neuro-Symbolic Integration Brings Causal and Reliable Reasoning Proofs
Sen Yang ${}^{1~{}2~{}3}$       Xin Li ${}^{2~{}3}$        Leyang Cui       Lidong Bing ${}^{2~{}3}$       Wai Lam 1 1 The Chinese University of Hong Kong 2 DAMO Academy, Alibaba Group 3 Hupan Lab, 310023, Hangzhou, China senyang.stu@gmail.com     {xinting.lx, l.bing}@alibaba-inc.com neal19951101@gmail.com     wlam@se.cuhk.edu.hk
@@ 1 Introduction
Overall, the contributions of this work are:
(1) proposing an alternative way for reasoning interpretability by making use of logs produced by declarative solvers;
(2) showcasing a Prolog-based implementation, which yields promising results regarding both final answers and intermediate steps.

### v3

# Sen Yang ${}^{1~{}2}$       Xin Li 2        Leyang Cui       Lidong Bing 3       Wai Lam 1

@@ Abstract
Two lines of approaches are adopted for complex reasoning with LLMs. One line of work prompts LLMs with various reasoning structures, while the structural outputs can be naturally regarded as intermediate reasoning steps. Another line of work adopt LLM-free declarative solvers to do the reasoning task, rendering higher reasoning accuracy but lacking interpretability due to the black-box nature of the solvers. Aiming to resolve the trade-off between answer accuracy and interpretability, we present a simple extension to the latter line of work. Specifically, we showcase that the intermediate search logs generated by Prolog interpreters can be accessed and interpreted into human-readable reasoning proofs. As long as LLMs correctly translate problem descriptions into Prolog representations, the corresponding reasoning proofs are ensured to be causal and reliable. On two logical reasoning and one arithmetic reasoning datasets, our framework obtains significant improvements in terms of both answer accuracy and reasoning proof accuracy. Our code is released at https://github.com/DAMO-NLP-SG/CaRing.
@@
Neuro-Symbolic Integration Brings Causal and Reliable Reasoning Proofs thanks:   The work described in this paper is partially supported by a grant from the Research Grant Council of the Hong Kong Special Administrative Region, China (Project Code: 14200620).  thanks:   This work was supported by Alibaba Group through the Alibaba Innovative Research (AIR) Program (TA2217728).
Sen Yang ${}^{1~{}2}$       Xin Li 2 thanks:   XL is the corresponding author.       Leyang Cui       Lidong Bing 3       Wai Lam 1 1 The Chinese University of Hong Kong 2 DAMO Academy, Alibaba Group 3 Shanda AI Research Institute senyang.stu@gmail.com     xinting.lx@alibaba-inc.com lidong.bing@shanda.com        neal19951101@gmail.com wlam@se.cuhk.edu.hk